# Polish Academy of Sciences

### v2

@@ Abstract
The article presents an innovative approach to the visualization of multidimensional data, using icons inspired by Chernoff faces. The approach merges classical projection techniques with the assignment of particular data dimensions to mimic features, capitalizing on the natural ability of the human brain to interpret facial expressions. We introduce a semantic division of data dimensions into intuitive and technical categories, assigning the former to avatar features and projecting the latter into a hyperspace of four, or potentially more dimensions. The technique is implemented as a plugin to the dpVision open-source image handling platform. The plugin allows the data to be interactively explored in the form of a swarm of avatars whose position in hyperspace as well as facial features represent various aspects of the data. Sample visualizations, based on synthetic test data as well as the 12-dimensional database on Portuguese Vinho Verde wines, confirm the usefulness of our approach to the analysis of complex data structures.
@@ 1 Introduction
Multidimensional data sets are a very rich resource in many applications, constantly gaining importance as both computer hardware and algorithms are increasingly capable of handling them. One element which is struggling to keep up with this development is the human visual system (and human senses in general), naturally built for the 3D space, and lacking the ability to perceive, or even imagine, a space of a higher dimensionality.
There are two ways to cope with the problem:
  1. 1.
    reduce the dimensionality of the data by taking a 3D projection or cross-section (possibly a thick-layer cross-section) of the data;
  2. 2.
    increase the dimensionality of the display by bestowing the representations of the points with additional characteristics besides their coordinates. Those additional properties can include color, size, shape parameters, any quantifiable visual peculiarities. Each data point is then represented by a more elaborate token, which in this text will be referred to as an avatar. One well-known approach is the use of Chernoff faces [1], icons resembling a human face, with data values converted into parameters such as rounded/oblong shape of the face, hair color and length, smiling/frowning curvature of lips etc. One advantage of Chernoff faces is that they tap into the natural ability of the human observer to identify and interpret human facial traits and expressions. However, this capacity also introduces challenges, as human observers may be influenced by the emotional connotations carried by various types and expressions of human or human-like faces. Perceiving an avatar as attractive/ugly or hostile/friendly adds a factor which is not necessarily related to the actual data and can seriously distort the analysis. To overcome this problem, Lardelli [2] proposed alternative glyphs in the form of trees. Trees are very familiar to humans, but seldom evoke strong emotions.
@@
In this work, we propose a novel approach that semantically separates data dimensions into two groups: those that are intuitively meaningful to the user (e.g., sensory, categorical, interpretable parameters), and those that are abstract or technical. Users may assign meaningful dimensions to either geometric space or avatar visual features. If fewer than the required number of dimensions are assigned, the system automatically supplements them using anonymous dimensions, which are processed via PCA to generate missing spatial or visual features. This approach is implemented as an interactive plugin for the dpVision platform [3, 4], allowing users to explore a cloud of 3D avatars positioned in a 4D data space and visualized through dynamic projections into 3D space (see Figure 1). By merging spatial navigation with avatar-based representation, our plugin provides a human-centered and intuitive visualization of complex multidimensional data.
@@
To validate the proposed approach, we use two types of multivariate datasets: a) an artificial 4D structure designed for easy geometric interpretation, and b) the publicly available dataset containing 12-dimensional data on Portuguese Vinho Verde wines [5, 6].
2 STATE OF THE ART
Many approaches have been tested to make multivariate point clouds visible to the human observer. The human senses of vision and proprioception, and, to a lesser extent, hearing, have developed to deal with the natural 3D space, and our imagination is also strongly limited by it; the ubiquitous use of 2D screens and printed material further limits the resources at our disposal. Consequently, authors have contrived many ingenious ways of projecting (in a very broad sense of the word) multidimensional realities into a 3D or 2D visual space.
The parallel coordinate plot [7] reduces the n-dimensional coordinate system to the 2D display plane by placing the coordinate axes in parallel (usually in a vertical direction) and representing each nD data point as a polyline with individual coordinate values as its nodes (Figure 2 (b, c)). An early application of this format in medicine was known as liver profiles, representing a single nD (typically, 6D) point as a polyline or as a set of parallel vertical bars of various height (Figure 2 (a)). When multiple nD points need to be displayed, the polylines are colored to keep them distinguishable (Figure 2 (b)). A smoothed version of this construct is the Andrews plot (Figure 2 (c)) [8].
A similar approach in a different (radial) coordinate system uses radial coordinates (Figure 2 (d)). While carrying the same information as parallel coordinates, its advantage is a somewhat easier visual comparison of plots, especially in relatively low dimensionalities (up to about 10).
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 2: Parallel and radial axis formats
A trellis chart (Figure 3) is a compact array of small 2D plots (usually all of the same type), showing relationships between two dimensions at a time, allowing the user or reader to examine them in a more comprehensive way than if they were spread over several pages or screens. A classic setup is a full set of all $\binom{N}{2}=\frac{n(n-1)}{2}$ possible pairs of dimensions, fitted into one page or screen.
Figure 3: A trellis chart
Purdom [9] provides a review of multivariate presentation, going through the four types of variables: nominal (categorical), ordinal, discrete, and continuous. Her paper includes large format trellis charts, high-resolution maps using color for a third dimension, and the use of Principal Component Analysis (PCA) to choose linear combinations of variables as aggregate "dimensions" representing the most pertinent aspect of the data.
Figure 4: 4D hyperbox
A more condensed representation was introduced by B. Alpern and L. Carter [10]. They represent the axes of an n-dimensional coordinate system as a 2D pencil of line segments emanating from a common origin and contained in an angle of less than 180 degrees. The $n-1$ additional copies of each segment are used to form a mesh of $\binom{N}{2}$ parallelograms, representing the coordinate planes of the hyperspace in a hyperbox (Figure 4), displayed with the physical appearance of a multifaceted 3D convex polyhedron in perspective.
Despite the large number of methods proposed for the visualization of multidimensional data, most of them focus either on mathematical reduction of dimensionality (such as PCA or t-SNE), or on geometric representations like parallel coordinates, radial plots, or static glyphs. These methods typically treat all data dimensions equally, without considering how understandable each dimension is to the human observer.
@@ 2 STATE OF THE ART
An alternative, cognitively grounded approach was proposed by Pflughoeft et al. [11], who introduced data avatars — metaphorical figures designed to visually encode multidimensional data in ways that align with human intuitive perception. Their work, based on dual-process cognitive theory, shows that representations tailored to intuitive understanding can improve user performance in data interpretation tasks.
We propose an approach that introduces two complementary elements not present in previous work. First, we clearly separate the dimensions of the dataset into two groups: those that are intuitively meaningful to users (such as quality ratings or sensory features), and those that are more abstract or technical (such as chemical coefficients). The intuitive features are assigned to visual characteristics of avatars—inspired by Chernoff faces—while the abstract features are reduced or projected into 3D spatial coordinates.
Second, we combine this concept with interactive 3D visualization, allowing users to explore a swarm of avatars from different angles and positions. While face-based visualization is known from Chernoff’s work, it has mostly been used in two-dimensional form. We are not aware of any previous system that integrates avatar-based glyphs with interactive spatial navigation and dimensional filtering in a 3D scene.
3 NOTATION AND TERMINOLOGY
@@
Throughout this paper, we use consistent notation to describe data dimensions, transformation matrices, and visualization parameters. The key symbols and abbreviations are summarized in Table 1, and are referenced throughout the description of the projection procedure and implementation details.
@@
$V$ View matrix $V\in\mathbb{R}^{(k+1)\times(k+1)}$
@@
4 PARADIGM
Multidimensional data sets often contain a mixture of parameters: some of them intuitively understandable to humans, and others more abstract or technical. To maximize interpretability, we propose a semantic division of data dimensions into two groups:
  • Intuitive dimensions, such as quality ratings or sensory properties, are mapped onto visual characteristics of avatars (e.g., face shape, expressions, color).
  • Technical dimensions, such as chemical or structural parameters, are projected into spatial coordinates in a 3D observation space after optional dimensionality reduction.
@@ 4 PARADIGM
A filtering matrix $F\in\mathbb{R}^{(k+m)\times n}$ is constructed to separate and process the dimensions, where $k$ is the number of spatial dimensions and $m$ the number of visual avatar features. The filtered data are:
$X_{\text{filtered}}=F\cdot X$
Here, each row of $F$ defines a new output feature (spatial or visual) as a linear combination of the original dimensions of $X$.
Spatial coordinates are obtained from a subset $F_{\text{spatial}}$, where $F_{\text{spatial}}\in\mathbb{R}^{k\times n}$, resulting in:
$X_{\text{spatial}}=F_{\text{spatial}}\cdot X$
Similarly, visual parameters are derived from $F_{\text{visual}}$, where $F_{\text{visual}}\in\mathbb{R}^{m\times n}$
$X_{\text{visual}}=F_{\text{visual}}\cdot X$
The spatial data are further transformed using a viewing matrix $V\in\mathbb{R}^{k\times n}$ (where typically $k=3$ or $4$):
$X_{\text{view}}=V\cdot X_{\text{spatial}}$
This dual mapping—avatars for intuitive features, spatial positioning for abstract features—provides a flexible and human-centered way to visualize high-dimensional data.
The conceptual framework of the proposed method, including the separation of intuitive and technical dimensions and the application of the filtering and viewing matrices, is illustrated in Figure 5.
5 IMPLEMENTATION
Based on the conceptual framework presented in the previous section, the practical implementation of the proposed visualization method involves the construction of appropriate transformation matrices, dimensionality reduction where necessary, and the assignment of selected data dimensions to spatial coordinates or avatar visual features.
@@ 5 IMPLEMENTATION
5.1 Assigning dimensions to categories
Note: To avoid confusion, we refer to dimensions when describing input data, and to features when referring to the visualized output.
@@ 5.1 Assigning dimensions to categories
5.2 Filtering matrix
The first step is to create the filtering matrix, which in the implementation has the following form:
@@ 5.2 Filtering matrix
To simplify later operations, we assume that the first $k$ rows of this matrix correspond to spatial features (in our implementation, these are the 4 features: X, Y, Z, T), and the following $m$ rows correspond to visual features (in our case: 10 features representing details of the avatar’s appearance). The matrix $F$ is initially filled with zeros and then completed based on the feature assignments:
  • First, for each of the $h_{S}$ spatial features and $h_{V}$ visual features selected by the user, 1 is set at the intersection of the row for that feature and the column of the original input dimension.
@@
5.3 View matrix
The implementation supports navigation in a $k$-dimensional space (typically 4D) using affine transformations. These transformations are represented by a square matrix of size $(k+1)\times(k+1)$, which combines both rotation and translation.
The core of this matrix is a $k\times k$ rotation matrix, generated by specifying the rotation plane (two coordinate axes) and the rotation angle.
@@ 5.3 View matrix
This allows smooth 4D navigation, including both rotation and panning of the data in the observation space.
5.4 Slab-based Filtering
To limit the number of visible data, a slab-based filtering mechanism is used. The goal is to retain only those data points that lie close to the viewing space—within a given distance slab_thickness from a hyperplane defined by the current rotation.

### v3

# Visualization of a multidimensional point cloud as a 3D swarm of avatars

@@ Abstract
This paper proposes an innovative technique for representing multidimensional datasets using icons inspired by Chernoff faces. Our approach combines classical projection techniques with the explicit assignment of selected data dimensions to avatar (facial) features, leveraging the innate human ability to interpret facial traits. We introduce a semantic division of data dimensions into intuitive and technical categories, assigning the former to avatar features and projecting the latter into a four-dimensional (or higher) spatial embedding. The technique is implemented as a plugin for the open-source dpVision visualization platform, enabling users to interactively explore data in the form of a swarm of avatars whose spatial positions and visual features jointly encode various aspects of the dataset. Experimental results with synthetic test data and a 12-dimensional dataset of Portuguese Vinho Verde wines demonstrate that the proposed method enhances interpretability and facilitates the analysis of complex data structures.
@@ 1 Introduction
Multidimensional datasets are a very rich resource in many applications, constantly gaining importance as both computer hardware and algorithms are increasingly capable of handling them. One element that is struggling to keep up with this development is the human visual system (and human senses in general), which is naturally built for 3D space and cannot perceive or even imagine a space of higher dimensionality.
There are two ways to cope with the problem: (i) reduce the dimensionality of the data by taking a 3D projection or cross-section (possibly a thick-layer cross-section) of the data; (ii) increase the dimensionality of the display by bestowing the representations of the points with additional characteristics besides their coordinates. Those additional properties can include color, size, shape parameters, and any quantifiable visual peculiarities. Each data point is then represented by a more elaborate token, which in this text will be referred to as an avatar.
One well-known approach to avatar-based representation is the use of Chernoff faces [1], icons resembling a human face, where data values are converted into parameters such as the rounded or oblong shape of the face, hair color, and length, smiling or frowning curvature of lips, etc. Chernoff faces leverage the natural ability of the human observer to identify and interpret facial traits and expressions and have been widely used for intuitive visualization of multivariate data.
However, Chernoff faces present several limitations, especially in the context of complex, high-dimensional, or interactive data analysis. The number of distinct visual features that can be reliably mapped onto a face is inherently limited, constraining the dimensionality that can be visualized. Most implementations are static and two-dimensional, limiting opportunities for dynamic exploration or spatial context. Importantly, the human perception of faces is subject to emotional and subjective biases—for example, certain expressions may be perceived as ”friendly” or ”untrustworthy” regardless of the underlying data [2]. Furthermore, Chernoff faces typically treat all variables equally without distinguishing between dimensions that are intuitively meaningful to the user and those that are more abstract or technical. It can lead to loss of interpretability or visual clutter, especially as the number of variables increases. To overcome some of these issues, alternative glyphs have been proposed, such as trees [2], which are familiar to humans but evoke less emotion.
@@
In this work, we propose a novel approach that semantically separates data dimensions into two groups: those that are intuitively meaningful to the user (e.g., sensory, categorical, interpretable parameters), and those that are abstract or technical. Users may assign meaningful dimensions to either geometric space or avatar visual features. If fewer than the required number of dimensions are assigned, the system automatically supplements them using anonymous dimensions, which are processed via Principal Component Analysis (PCA) to generate missing spatial or visual features. This approach is implemented as an interactive plugin for the dpVision platform [3, 4], allowing users to explore a cloud of 3D avatars positioned in a 4D data space and visualized through dynamic projections into 3D space (see Figure 1). By merging spatial navigation with avatar-based representation, our plugin provides a human-centered and intuitive visualization of complex multidimensional data.
@@
To validate the proposed approach, we use two types of multivariate datasets: a) two synthetic datasets designed to illustrate interpretability and the effects of variable assignment—one modeling fictional politicians and another simulating a diverse set of soft drinks; and b) the publicly available dataset containing 12-dimensional data on Portuguese Vinho Verde wines [5, 6].
The remainder of this paper is organized as follows. Section 2 reviews existing methods for multidimensional data visualization and highlights their limitations. Section 3 introduces the conceptual framework underlying our approach, including the semantic division of data dimensions, and describes the technical implementation of the visualization system as a plugin for the dpVision platform. Section 4 presents experiments on both synthetic and real-world datasets and explains the process of interactive exploration of the data cloud. Section 5 summarizes the main contributions of the work and outlines directions for future research. Finally, Section 6 concludes the paper.
2 Related work
The problem of visualizing multidimensional data has inspired a diverse set of techniques, each with its strengths and limitations. Projection-based methods, such as Principal Component Analysis (PCA) [7], t-distributed stochastic neighbor embedding (t-SNE) [8], and Uniform Manifold Approximation and Projection (UMAP) [9], are commonly used to reduce high-dimensional datasets to two or three dimensions, facilitating the discovery of patterns, clusters, or outliers.
Coordinate-based visualizations, such as parallel coordinates [10] and radial plots [11], allow each variable to be explicitly represented as an axis. Parallel coordinates, in particular, have been widely used in medicine (e.g., ”liver profiles”), and their variants, such as Andrews plots [12], provide additional smoothing and compactness. These approaches are effective for moderate numbers of dimensions but can become visually cluttered as dimensionality increases.
Matrix-based and small multiples techniques, including trellis charts [13], offer compact overviews of pairwise relationships between variables, enabling users to examine a large number of possible interactions.
Glyph-based approaches, most notably Chernoff faces [1], map multiple variables onto visual features of an icon or avatar, leveraging innate human abilities for face and pattern recognition. However, the subjectivity and emotional connotations of human-like glyphs can bias interpretation, and most glyph-based methods are static and limited to two dimensions.
Dynamic and interactive exploration techniques have also been developed, such as the grand tour and guided tour [14] and the slice tour [15]. These methods present users with sequences of projections or thin slices through the data space, sometimes augmented with glyphs for detailed inspection (e.g., the tourr package for R). While powerful for uncovering high-dimensional structures, these approaches typically treat all variables as equivalent, making them challenging for non-experts to interpret and understand.
@@ 2 Related work
More recently, cognitively grounded approaches have been explored. Pflughoeft et al. [16] introduced ”data avatars”—metaphorical glyphs designed to visually encode multidimensional data in ways that align with intuitive human perception. Their work, based on dual-process cognitive theory, demonstrated that such designs could improve interpretation; however, they did not address interactive navigation or explicit semantic division of variables.
Compared to the above methods, our approach introduces two key innovations. First, we enable a semantic split of the dataset into two groups: dimensions that are intuitively meaningful to users (such as sensory or quality ratings) and those that are more technical or abstract (such as chemical or process parameters). The intuitive features are mapped to avatar visual traits (inspired by Chernoff faces), while the technical features are projected into spatial coordinates, optionally using dimensionality reduction (e.g., PCA). Second, we integrate these glyphs into an interactive 3D environment, allowing users to explore a ”swarm” of avatars, dynamically rotate the viewpoint, and apply slab-based filtering to reveal local structure.
Our slab filtering is conceptually related to the slice tour [15], as both select points near a hyperplane for visualization. However, while the slice tour typically focuses on very thin slices (showing only points lying almost exactly on a hyperplane), our approach uses a slab of configurable thickness, retaining all points within a certain distance from the projection plane. This design supports the practical exploration of moderately sized datasets, reducing visual clutter while still revealing structural patterns. To our knowledge, this is the first system to combine avatar-based glyphs with interactive 3D navigation, explicit semantic variable assignment, and flexible local filtering in this manner.
3 Methodology
@@
Throughout this paper, we use consistent notation to describe data dimensions, transformation matrices, and visualization parameters. The key symbols and abbreviations are summarized in Table 1 and are referenced throughout the description of the projection procedure and implementation details.
@@
$V$ View matrix: for spatial rotation in $k$-dimensional space ($k\times k$), or full affine transformation with translation in homogeneous coordinates ($(k+1)\times(k+1)$) $V\in\mathbb{R}^{k\times k}$ or $V\in\mathbb{R}^{(k+1)\times(k+1)}$
@@
3.1 Paradigm and semantic assignment of variables
In practical applications, multidimensional datasets often contain both variables that are intuitively meaningful to users (such as sensory ratings and categorical or interpretable parameters) and features that are more technical or abstract (such as chemical coefficients or structural process data). To maximize interpretability, our approach introduces a semantic division of data dimensions into two groups:
  • Intuitive dimensions—readily interpreted by users—are mapped onto visual characteristics of avatars (e.g., face shape, expressions, color).
  • Technical dimensions—more abstract or difficult to interpret—are projected into spatial coordinates in the observation space, possibly after dimensionality reduction (e.g., PCA).
The assignment of variables to these groups should be guided by domain knowledge and the analysis’s objectives. Features that are critical for individual profiling or human interpretation should be mapped to avatar features, while those essential for discovering structures or clusters may be included in spatial coordinates. Redundancy should be avoided (e.g., highly correlated features mapped to multiple avatar traits), and important variables should not be left out of both avatar and spatial assignment, as this could diminish interpretability.
@@ 3.1 Paradigm and semantic assignment of variables
To implement the semantic assignment of variables, we construct a filtering matrix $F\in\mathbb{R}^{(k+m)\times n}$ that separates and processes the dimensions, where $k$ is the number of spatial dimensions and $m$ is the number of avatar visual features. The filtered data are given by Eq. (1):
$X_{\text{filtered}}=F\cdot X.$ (1)
Each row of $F$ defines a new output feature (spatial or visual) as a linear combination of the original dimensions of $X$.
Spatial coordinates are obtained by selecting the appropriate subset $F_{\text{spatial}}\in\mathbb{R}^{k\times n}$, as shown in Eq. (2):
$X_{\text{spatial}}=F_{\text{spatial}}\cdot X,$ (2)
while visual parameters are derived from $F_{\text{visual}}\in\mathbb{R}^{m\times n}$, according to Eq. (3):
$X_{\text{visual}}=F_{\text{visual}}\cdot X.$ (3)
Finally, the spatial data are further transformed for display using a view matrix $V$, as in Eq. (4):
$X_{\text{view}}=V\cdot X_{\text{spatial}}$ (4)
where $V$ is typically a $k\times k$ rotation matrix. In practical applications, particularly for interactive 3D or 4D exploration, we use an affine $(k+1)\times(k+1)$ matrix acting on homogeneous coordinates, combining both rotation and translation (see Implementation).
This dual mapping—avatars for intuitive features, spatial positioning for technical features—provides a flexible and human-centered way to visualize high-dimensional data (see Figure 2).
Guidelines and limitations. The semantic assignment should be made carefully, considering both user needs and analytical goals. Important features should not be omitted or relegated solely to PCA-based projection, as their interpretability may be lost (principal components represent blends of original variables). The division between intuitive and technical features is context-dependent and may vary by task or user group. We recommend iterative testing of assignments, consultation with domain experts, and documenting mapping choices. When PCA is used to process unassigned (anonymous) variables, users should inspect principal component loadings to understand which original features contribute most strongly to the resulting spatial dimensions. Although our current tool does not include interactive visualization of these loadings, such analysis can be performed using standard statistical software.
This framework strikes a balance between the interpretability of human-centered visualization and the structural insight provided by dimensionality reduction, remaining transparent about the trade-offs involved.
3.2 Implementation
Based on the conceptual framework presented in the previous section, the practical implementation of the proposed visualization method involves constructing appropriate transformation matrices, performing dimensionality reduction where necessary, and assigning selected data dimensions to spatial coordinates or avatar visual features.
@@ 3.2 Implementation
3.2.1 Assigning dimensions to categories
Note: To avoid confusion, we refer to dimensions when describing input data and to features when referring to the visualized output.
@@ 3.2.1 Assigning dimensions to categories
3.2.2 Filtering matrix
The first step, corresponding to Eq. (1), is to create the filtering matrix, which in the implementation has the following form:
@@ 3.2.2 Filtering matrix
To simplify later operations, we assume that the first $k$ rows of this matrix correspond to spatial features (in our implementation: X, Y, Z, T), and the following $m$ rows correspond to visual features (e.g., avatar appearance details). This separation is formally described by Eqs. (2) and (3). The matrix $F$ is initially filled with zeros and then completed based on the feature assignments:
  • First, for each of the $h_{S}$ spatial features and $h_{V}$ visual features selected by the user, one is set at the intersection of the row for that feature and the column of the original input dimension.
@@
3.2.3 View matrix
The view matrix $V$ is implemented as an affine $(k+1)\times(k+1)$ matrix, combining spatial rotation (the $k\times k$ upper-left block) and translation (the final column) for interactive navigation in homogeneous coordinates. This approach enables efficient and flexible real-time manipulation of the data cloud in 3D or 4D space.
@@ 3.2.3 View matrix
This corresponds directly to the transformation defined in Eq. (4), where the spatial coordinates are mapped into the observation space for visualization. This procedure enables smooth 4D navigation, including both rotation and panning of the data cloud in the observation space.
3.2.4 Slab-based Filtering
To limit the number of visible data, a slab-based filtering mechanism is employed. The goal is to retain only those data points that lie close to the viewing space—within a given distance slab_thickness from a hyperplane defined by the current rotation.